A 30-m3 tank is filled with 3480 moles of an ideal gas. The temperature of the gas is initially 305 K, but it is raised to 750 K.

1)What is the change in the pressure of the gas inside the tank?Pa

2)Let's assume that the tank is made out of a material which has a coefficient of linear expansion of 6.5 × 10-5 K-1, and that the temperature of the tank is always the same as the temperature of the gas. Remember that the space inside the tank will expand as if it were made of the material surrounding it. What is the change in the volume of the tank due to the increase in temperature from 305 K to 750 K?m3

3)Given that the tank expands a bit, what is the actual change in the gas pressure due to the increase in temperature?
Answer #1


For id eal eas P&T Pin Tin Ts 2.459 AM 305 3480 8.314x305 3o nRTin 294149- 32 Pa AP P-P 2.459 P honge in pressun AP=1.459 P 1 30+2-6 final volume - 32-6 m nRTs 11 3480x8. 314x 75o 32.6 665630. o6 Pa change in Pressure AP= P-P = 665630-06-294149-32 AP